Ik NG 12 ft El -- 2k 12 -- |--1201 - Fig. P4.9Find all the bar forces
Answer #1

Zero force members; - EF, so Ferro also Al is an axial force member, 4 reaction EMCO CD FACX12 = 2x12 + 2x24+1X36-2812 Fac=7kplease include a horizontal reaction at B inaFBD of joint B in last figure. Rbx = 2+2+1 = 5k towards left.